# Understanding and Mitigating Errors of LLM-Generated RTL Code

### v1

@@ Abstract
Despite the promising potential of large language model (LLM) based register-transfer-level (RTL) code generation, the overall success rate remains unsatisfactory. Errors arise from various factors, with limited understanding of specific failure causes hindering improvement. To address this, we conduct a comprehensive error analysis and manual categorization. Our findings reveal that most errors stem not from LLM reasoning limitations, but from insufficient RTL programming knowledge, poor understanding of circuit concepts, ambiguous design descriptions, or misinterpretation of complex multimodal inputs. Leveraging in-context learning, we propose targeted error correction techniques. Specifically, we construct a domain-specific knowledge base and employ retrieval-augmented generation (RAG) to supply necessary RTL knowledge. To mitigate ambiguity errors, we introduce design description rules and implement a rule-checking mechanism. For multimodal misinterpretation, we integrate external tools to convert inputs into LLM-compatible meta-formats. For remaining errors, we adopt an iterative debugging loop (simulation-error localization-correction). Integrating these techniques into an LLM-based framework significantly improves performance. We incorporate these error correction techniques into a foundational LLM-based RTL code generation framework, resulting in significantly improved performance. Experimental results show that our enhanced framework achieves 91.0% accuracy on the VerilogEval benchmark, surpassing the baseline code generation approach by 32.7%, demonstrating the effectiveness of our methods.
Index Terms:
Large Language model, RTL code generation, RTL error correction
@@ I Introduction
Recent advances in Large Language Models (LLMs) [1] have significantly reshaped the landscape of automated code generation. Specialized code generation models such as Codex[2], CodeGen[3] and foundational LLMs such as GPT[4] and DeepSeek[5] have demonstrated impressive capabilities in producing syntactically correct and functionally meaningful code across a wide range of programming languages and tasks. However, while LLMs achieve high success rates for mainstream programming languages like Python, their performance on hardware description languages (HDLs)[6] such as Verilog remains noticeably lower. This discrepancy is largely due to the scarcity of high-quality Verilog code in the pretraining corpora, as well as the inherent domain-specific complexity of RTL designs in general.
@@
Recent research has taken important steps toward addressing these RTL code generation challenges. Approaches such as VeriGen [8] ,VerilogEval [9], RTLCoder [10], Betterv [11] and DeepRTL[12] improve LLM performance by fine-tuning on larger domain-specific Verilog datasets. Other methods, including RTLFixer [13] and HDLDebugger [14], adopt retrieval-augmented generation (RAG) strategies, incorporating external knowledge bases or repair examples to compensate for the models’ limitations in hardware design and debugging. Recognizing the difficulty of resolving errors in a single step, recent systems such as AutoChip [15], MEIC [16], and VeriAssistant [17] employ iterative and multi-agent techniques to facilitate multi-round refinement and repair of LLM-generated Verilog code. Despite the varied methodologies, these approaches consistently demonstrate improved RTL code quality, suggesting that different techniques such as fine-tuning [18] can effectively address similar errors. However, these advancements often lack a systematic analysis of the root causes of the errors, raising the question of whether such errors are truly resolved or merely happen to be corrected by chance. This gap hinders further improvement in LLM-based RTL code generation.
In this work, we present a comprehensive analysis of errors in LLM-generated RTL code, categorizing them manually according to their underlying causes. Our findings indicate that the majority of errors do not stem from the reasoning limitations of LLMs, but rather from a lack of RTL programming knowledge, inadequate understanding of circuit design principles, ambiguous design specifications, or misinterpretation of complex multimodal inputs. Guided by these insights, we leverage the in-context learning capabilities of foundation models and introduce a set of targeted error correction techniques. Specifically, we construct a domain-specific knowledge base and apply retrieval-augmented generation (RAG) to provide LLMs with the specialized RTL programming and circuit design knowledge required for precise interpretation of design specification and accurate RTL code generation. To address issues arising from ambiguous design descriptions, we define a set of specification rules and implement a rule-checking mechanism to ensure clarity and completeness. For errors caused by misinterpretation of multimodal data, we integrate external tools to convert such inputs into meta-formats better suited for LLM processing. Additionally, to handle residual errors such as those resulting from incomplete contextual understanding in case of longer context, we adopt a classic iterative debugging loop consisting of simulation, error localization, and correction. We integrate all these techniques into a foundational LLM-based RTL code generation framework, achieving substantial improvements in both code correctness and overall generation quality.
@@
  • Based on the identified causes of errors, we propose a suite of correction techniques, including a domain-specific knowledge base with RAG, a rule-based mechanism for description refinement, multimodal input conversion tools, and an iterative simulation-based debugging loop.
  • We integrate the error correction mechanisms into a representative LLM-based RTL code generation framework, achieving 91.0% accuracy on the VerilogEval Benchmark —— surpassing a series of RTL generation framework.
  • We have open-sourced the erroneous RTL samples generated by GPT and other LLMs, along with their labeled results, analysis reports, and corresponding error-correction code on GitHub  [^1] https://github.com/zjz1222/RTLErrorAnalysis. We hope our work can contribute to advancing LLM-assisted hardware design.
@@ II-A LLMs for RTL Code Generation
With the rise of LLMs in the programming field, an increasing number of researchers are considering using LLMs for the automated generation of RTL designs from natural language descriptions. DAVE[19] emerged as an early study that fine-tuned GPT-2 [20] to automatically convert English specifications into Verilog code. Along this line, VeriGen[8] demonstrates that a fine-tuned open-source LLM (CodeGen-16B[3]) can outperform GPT-3.5 Turbo [21] in generating functionally correct Verilog code, highlighting the potential of specialized smaller models for hardware design automation. Subsequently, OriGen [22] enhances RTL dataset quality via code-to-code improvement, while AutoVCoder [23] utilizes data augmentation, two-phase fine-tuning, and domain-specific RAG. Betterv [11] fine-tunes LLMs on domain-specific data with generative discriminators for functionally correct code. DeepRTL [12] proposes a unified CodeT5+-based model, fine-tuned on Verilog code aligned with multi-level natural language descriptions. Following this, methods like Aivril[24], PromptV[25], and VerilogCoder[26] attempt to implement RTL generation using multi-agent approach. VRank[27] and MAGE[28] also employ multi-candidate sampling method to assist high-quality RTL generation. PREFACE[29] introduces a reinforcement learning based framework that enhances LLMs in generating formally verifiable code by iteratively refining prompts based on verification feedback.
@@ II-B LLMs for RTL Code Debugging
Despite the promising RTL code generation capabilities of LLMs, the success rate of the LLM-based RTL code generation remains unsatisfactory. An intuitive approach is to repeat the code generation procedures multiple times and select the best for the code generation, but this approach generally poses limited improvement and there is no guarantee for better results due to the uncertainty of LLM-based generation. Hence, debugging becomes a critical procedure to the LLM-based RTL code generation, and tremendous efforts have been devoted to fix the bugs and enhance the generated results continuously. For example, RTLFixer [13] combines the ReAct framework and RAG technology to fix compilation errors using a knowledge base of expert insights. HDLDebugger [14] enhances LLM error correction by integrating RAG with fine-tuning and a code sample library. AutoChip [15] employs a multi-round iterative approach guided by compilation and simulation feedback, while MEIC[16] employs a multi-agent framework with iterative refinement to correct both syntactic and functional errors in Verilog code. VeriAssistant[17] implements Verilog code repair through iterative LLM self-verification and self-correction. VeriDebug[30] introduces a contrastive representation and guided correction approach for automated Verilog debugging.
@@ III Error Analysis of LLM-Generated RTL Code
Figure 1: The basic prompt used for RTL code generation.
To understand the RTL code generation capabilities of foundational LLMs, we analyze the errors in the RTL code produced by these models and investigate their root causes. In our study, we use GPT models as representative LLMs and adopt VerilogEval as the benchmark suite for generating RTL code from natural language specifications. A standard RTL code generation prompt is shown in Fig. 1. The generated RTL code is verified and simulated using Icarus Verilog. Considering that models with different parameter scales and types vary in their ability to generate RTL code, to ensure the completeness of the analysis, we conducted experiments with the Qwen-Coder-32B-Instruction model which has a small scale and was additionally pre-trained on code, the general-purpose GPT-3.5 Turbo model of ordinary size, and the more capable GPT-4 Turbo model. And we respectively collected 83, 89, and 65 faulty programming samples. We manually examine the failed cases and classify them into two primary categories as listed below.
@@
TABLE I: Analysis of errors induced by insufficient knowledge of specialized RTL programming
Error Category Error Examples Golden Generation QwenCoder Samples GPT-3.5 Samples GPT-4 Samples
@@
Wire in Always Block output out; always @(*) begin ... case (state) ... B: out = 1; end output out; assign out = (state == B); 9 8 1
@@
Numerical Processing Logic Error // Count the number of 1s in the first two bits assign count1 = in[1:0]; // Count the number of 1s in the last bit ... assign count2 = count1 + in[2]; assign out = count2; assign out = in[0]+in[1]+in[2]; 2 1 0
@@
Vector Bit Selection Error module TopModule (input clk, ... output [3:1] g ); assign g[1] = (state == B); assign g[0] = 1’b0; module TopModule (input clk, ... output [3:1] g ); assign g[1] = (state == B); 1 3 0
@@
Inversion of Vector Slice Selection module TopModule ( input [99:0] in, ... output [99:1] out_any, ); assign out_any = in[99:1] | in[0:98]; module TopModule ( input [99:0] in, ... output [99:1] out_any, ); assign out_any = in[99:1] | in; 1 1 0
Incomplete RTL Code assign out = (sel == 0) ? in[0]: (sel == 1) ? in[1]: //Continue this pattern for all 256 inputs in[255]; assign out = in[sel]; 0 4 3
Variable Redefinition module TopModule(input clk, ..., output reg [7:0] q); reg [7:0] q; module TopModule(input clk, ..., output reg [7:0] q); 0 4 0
@@
Using Undefined Variables module TopModule (input c, d, output [3:0] mux_in ); wire w1, w2 ...; assign mux_in[0] = a & b ? w1 : w2; //The variables a and b are undefined. module TopModule (input c, d, output [3:0] mux_in ); assign mux_in[0] = c | d; 0 1 1
@@
Incorrect Use of Generate Statement module TopModule(input clk, ..., output reg [511:0] q); always @(posedge clk) begin ... genvar i; generate for (i = 1; i < 511; i = i + 1) begin: gen_rule90 always @(posedge clk) q[i] <= q[i-1] ^ q[i+1]; end endgenerate ... end module TopModule(input clk, ..., output reg [511:0] q); always @(posedge clk) begin integer i; for (i = 1; i < 511; i = i + 1) begin q[i] <= (q[i-1] ^ q[i+1]); end end 0 1 1
Mixing Blocking and Non-blocking Assignments always @(posedge clk) begin ... q[0] = (q[1] ^ 0); for (i = 1; i < 511; i = i + 1) q[i] <= (q[i-1] ^ q[i+1]); end always @(posedge clk) begin ... q[0] <= (q[1] ^ 0); for (i = 1; i < 511; i = i + 1) q[i] <= (q[i-1] ^ q[i+1]); end 0 1 0
@@
For different categories of LLMs, we conducted a detailed statistical analysis of the distribution of the two aforementioned error types, as shown in Fig 2. The results indicate that the majority of design errors (over 70%) stem from misinterpretations of design specifications, and this proportion increases with the model’s programming capability. In contrast, errors related to RTL-specific programming account for only a minor portion. A detailed analysis of both error categories is provided in the following subsections.
@@ III-A Insufficient Knowledge of Specialized RTL Programming
Although Verilog shares similarities with C-like languages, it incorporates specialized syntax and is deeply intertwined with circuit design concepts, making it substantially different from conventional high-level languages such as Python or C/C++. For example, Verilog enforces distinct assignment semantics for different contexts: blocking assignments and non-blocking assignments are used for combinational and sequential logic, respectively. Foundational LLMs often struggle with such distinctions due to the limited availability of Verilog-specific data in their training corpora. As a result, many RTL coding errors stem from insufficient understanding of RTL programming. According to our experiments on VerilogEval, we identified a total of 41 failed cases, which can be roughly categorized into four major groups, as shown in Table I.
Wire in Always Block: The LLM often fails to correctly interpret RTL variable usage specifications, leading to incorrect assignments to variables declared as wire types within always blocks, as shown in Table I. This type of error is more likely to occur in smaller-scale models such as Qwen-Coder-32B-Instruction. However, for larger-scale models like GPT-4 Turbo, as the training corpus and model parameters increase, the frequency of such errors decreases significantly.
Numerical Processing Logic Error: In some design contexts, LLMs are prone to errors in numerical processing logic. For instance, as illustrated in the Table  I. When implementing the function of counting the number of 1’s in the vector in[1:0], the LLM incorrectly equates in[1:0] with in[1] + in[0].
@@
Incomplete RTL Code: The LLM correctly interprets the design specifications, but generates only a portion of the required code. It often omits the remaining code inappropriately, particularly when the expected output is lengthy and repetitive. As shown in Table I, the model uses a comment to substitute the remaining branching cases in a typical switch block, instead of generating all 256 possible branches.
@@
Using Undefined Variables: In some RTL designs, LLMs may improperly reference undeclared variables during RTL generation. As shown in Table I, when implementing the mux_in[0] signal, the LLM incorrectly uses undefined variables ’a’ and ’b’, resulting in compilation errors.
Incorrect Use of Generate Statement: In some cases, LLMs may confuse ’generate for’ constructs with regular ’for’ statements. The case in Table I shows erroneous usage where the LLM attempted to iteratively update individual bits of register ’q’ through generate statements, violating synthesis rules and leading to malfunction.
Mixing Blocking and Non-blocking Assignments: In certain always blocks, LLMs exhibit a tendency to mix blocking and non-blocking assignments, resulting in functional errors.
Based on the statistics above, we can observe that Qwen-Code-32B-Instruction tends to make errors in numerical and vector processing logic, GPT-4 Turbo often generates incomplete code or misuses advanced RTL syntax like Generate, while GPT-3.5 Turbo produces a wider variety of errors. This indicates that both specialized training on code data for smaller models and simply scaling up model parameters and training data can help reduce errors caused by insufficient RTL programming knowledge in LLMs.

### v2

@@ Abstract
Despite the potential of large language model (LLM) based register-transfer-level (RTL) code generation, the overall success rate remains unsatisfactory, with limited understanding of the underlying causes of errors. To address this gap, we conduct a comprehensive analysis of errors in generated RTL code and manually categorize them based on their underlying causes. Our findings reveal that most errors stem not from the reasoning capabilities of LLMs, but rather from a lack of RTL programming knowledge, insufficient understanding of circuit concepts, ambiguous design descriptions, or misinterpretation of complex multimodal inputs. Building on this insight, we leverage the in-context learning capabilities of foundational LLMs and propose a set of error correction techniques to address these issues effectively. Specifically, we construct a domain-specific knowledge base and employ retrieval-augmented generation (RAG) to supply LLMs with the specialized knowledge required for RTL code generation. To mitigate errors arising from ambiguous descriptions, we introduce a set of design description rules and implement a rule-checking mechanism that helps clarify and complete the input specifications. For errors caused by misinterpretation of multimodal data, we integrate external tools that convert such inputs into meta-formats more compatible with LLM processing. For the remaining errors such as missing details of longer context, we adopt a classic iterative debugging loop i.e. simulation-error localization-error correction to gradually improve the generated RTL code. We incorporate these error correction techniques into a foundational LLM-based RTL code generation framework, resulting in significantly improved performance. Experimental results show that our enhanced framework achieves 98.1% accuracy on the VerilogEval benchmark based on Deepseek-v3.2-Speciale, demonstrating the effectiveness of our methods.
@@ I Introduction
Recent advances in Large Language Models (LLMs) [14] have significantly reshaped the landscape of automated code generation. Specialized code generation models such as Codex[2], CodeGen[15] and foundational LLMs such as GPT[1] and DeepSeek[8] have demonstrated impressive capabilities in producing syntactically correct and functionally meaningful code across a wide range of programming languages and tasks. However, while LLMs achieve high success rates for mainstream programming languages like Python, their performance on hardware description languages (HDLs) such as Verilog remains noticeably lower. This discrepancy is largely due to the scarcity of high-quality Verilog code in the pretraining corpora, as well as the inherent domain-specific complexity of RTL designs in general.
@@
Recent research has taken important steps toward addressing these RTL code generation challenges. Approaches such as VeriGen [22] ,VerilogEval [10], RTLCoder [12], CodeV-R1[31] improve LLM performance by fine-tuning and reinforcement learning on larger domain-specific Verilog datasets. Other methods, including RTLFixer [24] and HDLDebugger [28], adopt retrieval-augmented generation (RAG) strategies, incorporating external knowledge bases or repair examples to compensate for the models’ limitations in hardware design and debugging. Recognizing the difficulty of resolving errors in a single step, recent systems such as AutoChip [23], MEIC [27] and MAGE[29] employ iterative and multi-agent techniques to facilitate multi-round refinement and repair of LLM-generated Verilog code. Despite the varied methodologies, these approaches consistently demonstrate improved RTL code quality, suggesting that different techniques such as fine-tuning [11] and can effectively address similar errors. However, these advancements often lack a systematic analysis of the root causes of the errors, raising the question of whether such errors are truly resolved or merely happen to be corrected by chance. This gap hinders further improvement in LLM-based RTL code generation.
In this work, we present a comprehensive analysis of errors in LLM-generated RTL code, categorizing them manually according to their underlying causes. Our findings indicate that the majority of errors do not stem from the reasoning limitations of LLMs, but rather from a lack of RTL programming knowledge, inadequate ambiguous design specifications, misinterpretation of complex multimodal inputs or understanding of circuit design principles. Guided by these insights, we leverage the in-context learning capabilities of foundation models and introduce a set of targeted error correction techniques. Specifically, to address issues arising from ambiguous design descriptions, we define a set of specification rules and implement a rule-checking mechanism to ensure clarity and completeness. For errors caused by misinterpretation of multimodal data, we integrate external tools to convert such inputs into meta-formats better suited for LLM processing. Furthermore, we apply retrieval-augmented generation (RAG) to provide LLMs with the specialized RTL programming and circuit design knowledge required for precise interpretation of design specifications and accurate RTL code generation, supported by the construction of a domain-specific knowledge base. Additionally, to handle residual errors such as those resulting from incomplete contextual understanding in case of longer context, we adopt a classic iterative debugging loop consisting of simulation, error localization, and correction. We integrate all these techniques into a foundational LLM-based RTL code generation framework, achieving substantial improvements in both code correctness and overall generation quality.
@@
  • Based on the identified causes of errors, we propose a suite of correction techniques, including a rule-based mechanism for description refinement, multimodal input conversion tools, a domain-specific knowledge base with RAG, and an iterative simulation-based debugging loop.
  • We integrate the error correction mechanisms into a representative LLM-based RTL code generation framework, achieving 98.1% accuracy on the VerilogEval Benchmark based on Deepseek-v3.2-Speciale —— surpassing a series of RTL generation framework.
  • We have open-sourced the erroneous RTL samples generated by LLMs, along with their labeled results, analysis reports, and corresponding error-correction code on GitHub  [^1] https://github.com/zjz1222/RTLErrorAnalysis. We hope our work can contribute to advancing LLM-assisted hardware design.
@@ II-A LLMs for RTL Code Generation
With the rise of LLMs in the programming field, an increasing number of researchers are considering using LLMs for the automated generation of RTL designs from natural language descriptions. DAVE[18] emerged as an early study that fine-tuned GPT-2 [20] to automatically convert English specifications into Verilog code. Along this line, VeriGen[22] demonstrates that a fine-tuned open-source LLM (CodeGen-16B[15]) can outperform GPT-3.5 Turbo [16] in generating functionally correct Verilog code, highlighting the potential of specialized smaller models for hardware design automation. Subsequently, OriGen [4] enhances RTL dataset quality via code-to-code improvement. CodeV-R1 [31] proposes a RLVR framework for training Verilog-generating LLMs. ReasoningV [19] introduces a high-quality Verilog code dataset that incorporates reasoning. Following this, methods like Aivril[21], PromptV[13], and VerilogCoder[6] attempt to implement RTL generation using multi-agent approach. VRank[30] and MAGE[29] also employ multi-candidate sampling method to assist high-quality RTL generation. HLSPilot[26] also introduced an LLM-based framework for HLS generation.
@@ II-B LLMs for RTL Code Debugging
Despite the promising RTL code generation capabilities of LLMs, the success rate of the LLM-based RTL code generation remains unsatisfactory. An intuitive approach is to repeat the code generation procedures multiple times and select the best for the code generation, but this approach generally poses limited improvement and there is no guarantee for better results due to the uncertainty of LLM-based generation. Hence, debugging becomes a critical procedure to the LLM-based RTL code generation, and tremendous efforts have been devoted to fix the bugs and enhance the generated results continuously. For example, RTLFixer [24] combines the ReAct framework and RAG technology to fix compilation errors using a knowledge base of expert insights. HDLDebugger [28] enhances LLM error correction by integrating RAG with fine-tuning and a code sample library. AutoChip [23] employs a multi-round iterative approach guided by compilation and simulation feedback, while MEIC[27] employs a multi-agent framework with iterative refinement to correct both syntactic and functional errors in Verilog code. VeriDebug[25] introduces a contrastive representation and guided correction approach for automated Verilog debugging.
@@ III Error Analysis of LLM-Generated RTL Code
TABLE I: Statistics of faulty design instances for different models on the VerilogEval benchmark11footnotetext: The release date corresponds to the latest version of the model available as of the study.
Category Model Name Release Date111* Corresponding author. Faulty Design
General DeepSeek-V3.2-Speciale[9] 2025.12 17
GPT-4 Turbo[17] 2024.04 65
GPT-3.5 Turbo[16] 2024.01 89
Code-Specific Qwen2.5-Coder-32B-Instruct[7] 2024.11 83
RTL-Specific CodeV-R1-RL-Qwen-7B[31] 2025.06 52
To understand the RTL code generation capabilities of foundational LLMs, we analyze the errors in the RTL code produced by these models and investigate their root causes.
Based on the VerilogEval benchmark, we conduct a comprehensive evaluation of the current large language models’ capability to generate RTL code by employing a diverse set of models that vary in training dataset, parameter scale, and reasoning ability. The evaluation results are presented in Table I.
Our assessment strategy is tailored to the characteristics of the models. For standard text-completion models without reinforcement learning fine-tuning for reasoning tasks, such as GPT-3.5 Turbo, GPT-4 Turbo, and Qwen2.5-Coder-32B-Instruct, we adopt a direct Natural Language to Code (NL2Code) approach. For models specifically trained with reinforcement learning for reasoning tasks, like DeepSeek-V3.2-Speciale and CodeV-R1-RL-Qwen-7B, we utilize their recommended Reasoning configurations to fully leverage their code generation potential. The generated RTL code is verified and simulated using Icarus Verilog.
From the test results, a total of 306 faulty design instances are collected and analyzed. We manually examine the failed cases and classify them into two primary categories as listed below.
@@
TABLE II: Analysis of errors induced by insufficient knowledge of specialized RTL programming
Error Category Error Examples Golden Generation DS GPT-4 GPT-3.5 Qwen CodeV-R1
@@
Wire in Always Block output out; always @(*) begin ... case (state) ... B: out = 1; end output out; assign out = (state == B); 0 1 8 9 0
@@
Numerical Processing Logic Error // Count the number of 1s in the first ... assign count1 = in[1:0]; // Count the number of 1s in the last ... assign count2 = count1 + in[2]; assign out = count2; assign out = in[0]+in[1]+in[2]; 0 0 1 2 1
@@
Vector Bit Selection Error module TopModule (input clk, ... output [3:1] g ); assign g[1] = (state == B); assign g[0] = 1’b0; module TopModule (input clk, ... output [3:1] g ); assign g[1] = (state == B); 0 0 3 1 1
@@
Inversion of Vector Slice Selection module TopModule ( input [99:0] in, ... output [99:1] out_any, ); assign out_any = in[99:1] | in[0:98]; module TopModule ( input [99:0] in, ... output [99:1] out_any, ); assign out_any = in[99:1] | in; 1 0 1 1 0
Incomplete RTL Code assign out = (sel == 0) ? in[0]: (sel == 1) ? in[1]: //Continue this pattern for all 256 inputs in[255]; assign out = in[sel]; 0 3 4 0 0
Variable Redefinition module TopModule(input clk, ..., output reg [7:0] q); reg [7:0] q; module TopModule(input clk, ..., output reg [7:0] q); 0 0 4 0 0
@@
Using Undefined Variables module TopModule (input c, d, output [3:0] mux_in ); wire w1, w2 ...; assign mux_in[0] = a & b ? w1 : w2; //The variables a and b are undefined. module TopModule (input c, d, output [3:0] mux_in ); assign mux_in[0] = c | d; 0 1 1 0 0
@@
Incorrect Use of Generate Statement module TopModule(input clk, ..., output reg [511:0] q); always @(posedge clk) begin ... genvar i; generate for (i = 1; i < 511; i = i + 1) begin: gen_rule90 always @(posedge clk) q[i] <= q[i-1] ^ q[i+1]; end endgenerate ... end module TopModule(input clk, ..., output reg [511:0] q); always @(posedge clk) begin integer i; for (i = 1; i < 511; i = i + 1) begin q[i] <= (q[i-1] ^ q[i+1]); end end 1 1 1 0 1
@@
For different categories of LLMs, we conducted a detailed statistical analysis of the distribution of the different aforementioned error types, as shown in Fig 1. Increasing model scale or domain-specific RTL training can effectively reduce Type I errors. However, Type II errors — particularly insufficient understanding of circuit concepts — remain the primary bottleneck limiting LLMs’ ability to correctly generate RTL code.
@@ III-A Insufficient Knowledge of Specialized RTL Programming
Although Verilog shares similarities with C-like languages, it incorporates specialized syntax and is deeply intertwined with circuit design concepts, making it substantially different from conventional high-level languages such as Python or C/C++. For example, Verilog enforces distinct assignment semantics for different contexts: blocking assignments and non-blocking assignments are used for combinational and sequential logic, respectively. Foundational LLMs often struggle with such distinctions due to the limited availability of Verilog-specific data in their training corpora. As a result, many RTL coding errors stem from insufficient understanding of RTL programming. According to our experiments on VerilogEval, we identified a total of 41 failed cases, which can be roughly categorized into four major groups, as shown in Table II.
Wire in Always Block: The LLM often fails to correctly interpret RTL variable usage specifications, leading to incorrect assignments to variables declared as wire types within always blocks, as shown in Table II. This type of error is more likely to occur in smaller-scale models such as Qwen-Coder-32B-Instruction. However, for larger-scale models like GPT-4 Turbo, as the training corpus and model parameters increase, the frequency of such errors decreases significantly.
Numerical Processing Logic Error: In some design contexts, LLMs are prone to errors in numerical processing logic. For instance, as illustrated in the Table  II. When implementing the function of counting the number of 1’s in the vector in[1:0], the LLM incorrectly equates in[1:0] with in[1] + in[0].
@@
Incomplete RTL Code: The LLM correctly interprets the design specifications, but generates only a portion of the required code. It often omits the remaining code inappropriately, particularly when the expected output is lengthy and repetitive. As shown in Table II, the model uses a comment to substitute the remaining branching cases in a typical switch block, instead of generating all 256 possible branches.
@@
Using Undefined Variables: In some RTL designs, LLMs may improperly reference undeclared variables during RTL generation. As shown in Table II, when implementing the mux_in[0] signal, the LLM incorrectly uses undefined variables ’a’ and ’b’, resulting in compilation errors.
Incorrect Use of Generate Statement: In some cases, LLMs may confuse ’generate for’ constructs with regular ’for’ statements. The case in Table II shows erroneous usage where the LLM attempted to iteratively update individual bits of register ’q’ through generate statements, violating synthesis rules and leading to malfunction.
Based on the statistics presented above, distinct error patterns are identified across the models: Qwen2.5-Coder-32B-Instruct is prone to flawed logic in numerical and vector operations; GPT-4 Turbo tends to produce incomplete code or misuse advanced RTL syntax such as generate blocks; GPT-3.5 Turbo exhibits a wide variety of errors; and CodeV-R1-RL-Qwen-7B shows susceptibility to mistakes in specific numerical and vector processing designs.
We also observe that both large models fine-tuned with RL reasoning for programming tasks and smaller models trained specifically in the RTL domain contribute to mitigating the issue of Insufficient Knowledge of Specialized RTL Programming.